# Group-Control Motion Planning Framework for Microrobot Swarms in a Global Field

### v1

@@ 4.2 Control of Robot Orientation
We have showed that each group $G_{i}$ induces a vector field $g_{i}$ that just translates the robots in the group, and a vector field $h_{i}$ that just rotates the robots that are not in $G_{i}$. What we show next is that we can independently control the orientation of $m$ robots:
@@
The orientation of any $m$ robots can be controlled to any desired orientation $(\theta_{1},\ldots,\theta_{m})$.
Note that the rest of the swarm will rotate in place to some orientation that can not be directly controlled.
To illustrate this proposition, consider the rotation vector fields $g_{1},\ldots,g_{m-1}$ in (6) and the additional original rotation-only vector field $f_{m}$. All of these vector fields can rotate the robots without translating them. The proposition can then be proved by observing that using these vector fields, the orientations of the robots linearly depend on the arguments (rotation angles) of these vector fields. For example, for the 4-group case in Table 1, we get the equation:
$\begin{bmatrix}1&1&1&0&0&0\\ 1&0&0&1&1&0\\ 0&1&0&1&0&1\\ 1&1&1&1&1&1\end{bmatrix}^{T}*\begin{bmatrix}u_{1}&u_{2}&u_{3}&u_{4}\end{bmatrix}^{T}=q[2n+1:3n]-q[2n+1:3n](0)$ (7)
@@
It is easy to see that for $m$ groups, the rank of the matrix that appears in the equation above is $m$. This means that $m$ of the $n$ equations can be solved exactly for the desired inputs $u_{i}$.
@@ 4.3 Small-Time Local Controllability (STLC)
After we have identified the control vector fields in Eq. (6), we can proceed to show that the swarm is STLC in positions. Traditionally, we can do this by analyzing the closure of $\mathcal{F}$ under the Lie bracket operation. However, we will show that we can independently translate any robot in arbitrary direction without moving any other robot. This shows that the system is STLC in positions.
Suppose that we want to independently translate some robot $k\in\{1,\ldots,n\}$. Let $G_{\hat{k}}$ be any group that contains the robot $k$. Let $\operatorname{Rot}_{i_{1},\ldots,i_{m}}(\theta_{1},\ldots,\theta_{m})$ correspond to setting the orientation of robots $i_{1},\ldots,i_{m}$ to $\theta_{1},\ldots,\theta_{m}$ as described above. Also, assume that the members of the group $G_{\hat{k}}$ are robots $l_{1},\ldots,l_{j},k$.
@@
The control sequence $P_{1}(d)=(h_{\hat{k}}(d),\operatorname{Rot}_{l_{1},\ldots,l_{m}}(\pi,\ldots,\pi),h_{\hat{k}}(d))$ only translates robots $l_{m+1},\ldots,l_{j},k$.
@@
Observe that under the control sequence above, the first $m$ robots in $G_{\hat{k}}$ will travel back and forth while the rest of the robots in the group will undergo some translation.∎
Once we have the control sequence $P_{1}$, it is easy to see that a control sequence $(P_{1}(d),\operatorname{Rot}_{l_{m+1},\ldots,l_{2m}}(\pi,\ldots,\pi),P_{1}(d))$ will only translate robots $l_{2m+1},\ldots,l_{j},k$. By repeating this process we can therefore eliminate translation of all the robots $l_{1},\ldots,l_{j}$ except for the robot $k$. Of course, we can always rotate the robot $k$ appropriately so that the final translation is in the desired direction, showing that each robot can be translated independently, and that the system is STLC in positions.
@@ 5 Motion Planning
After we have shown that the swarm of MicroStressBots under group-control is STLC, we consider the motion planning problem for the swarm. Given our proof of STLC the problem is not how to move the swarm to a particular configuration, but how to move it efficiently. In other words, ideally we would like the robots to move to their desired configuration in parallel and not one by one. But planning such parallel motion directly is fraught with significant complexity due to the high dimension of the configuration space and the fact that we have a single control input. We have determined that effective motion planning for the global group control in high-dimensional state spaces should involve the design of motion primitives with subgroups.
The objective of designing motion primitives is to divide a robot group into smaller subgroups, each containing fewer robots. These primitives simplify the motion planning task by reducing the coupled motions in the group. We present primitives described by Lie bracket motions, as the higher the order of the Lie brackets, the fewer robots belong to it. It is important to note that, in this paper, these subgroups are logical constructs apart from $\log(n+2)+1$ number of physical groups. However, these logical groups can be physically manufactured to increase the number of physical groups, thereby enhancing control capabilities. Fig.1 shows the trade-offs between the complexity of control and planning. The order of Lie bracket $r_{k}$ is used to quantify the complexity. As we construct motion planning based on a higher order $r_{k}$, the control realizing $r_{k}$ order Lie brackets become more complex, while increasing the number of decoupled subgroups simplifies the planning (curve from left to right).
Figure 1: An illustration of control and motion planning complexity trade-offs. $r_{k}=1,r_{s},c_{1},c_{0}$ represent four cases: pure planning, pure control, controlling rotations, and decoupled subgroup, respectively.
In the following section, we discuss the complexity of motion planning through a formal abstraction of various levels of motion planning problems. Subsequently, we simulate the motion planning instantiations to demonstrate that the constructed subgroups can effectively reduce planning time.
@@ 5.1 Motion Planning Approximation Scheme
We start by giving some assumptions and definitions that are fundamental to the group-control robot motion planning problem. Several assumptions are provided below.
  • Robots are controlled by a global signal that operates them in groups, with one group activated at a time.
  • A predefined map of the area is available and known.
  • Each robot is assigned a starting point and a destination. The starting point includes the robot’s position and orientation, whereas the destination does not consider orientation.
@@
[…]
@@
Primitive Order The order of a primitive corresponding to the subgroup $p$ is the order of the Lie bracket used to realize it. The order of the Lie bracket is the number of generators from the set of vector fields $\mathcal{F}$ in Eq. (6) in it. In general, the higher the order of a Lie bracket, the fewer the number of robots move under that bracket. However, there is no one-to-one correspondence between the Lie bracket order and the number of robots it affects, as this also depends on the variety of rotation vector fields used to compute the Lie bracket. Fundamentally, the order of the Lie bracket indicates the difficulty of its implementation, while the number of robots affected by it reflects the complexity of the motion planning.
Definition 4
Motion Planning Problem Abstraction The motion planning problem for $n$ robots of level $r$, denoted as $\mathcal{M}_{s}^{g}(n,r)$, corresponds to the problem of finding a collision-free trajectory (sequence of motion primitives) for $n$ robots controlled by a global field from some initial state $s$ to a goal state $g$, where the motion primitives that are used have the primitive order at most $r$.
@@
Consider the motion planning problem $\mathcal{M}_{s}^{g}(n,r_{k})$ for $n$ robots, and $r_{k}\leq r_{s}$, where $r_{s}$ is the primitive order required for moving each individual robot (the highest order of the Lie bracket needed to prove STLC). If $r_{1}<r_{2}\leq r_{s}$, the complexity of the motion planning problem $\mathcal{M}_{s}^{g}(n,r_{2})$ is less than $\mathcal{M}_{s}^{g}(n,r_{1})$. Initially, when $r_{k}=r_{s}$, the complexity of the motion planning problem $\mathcal{M}_{s}^{g}(n,r_{s})$ is linear in robot number. Moreover, for any other $r_{k}$, the complexity of motion planning problem $\mathcal{M}_{s}^{g}(n,r_{k})$ has a loose upper bound $o(nL^{r_{s}-r_{k}})$, assuming moving along the first order Lie bracket has a constant complexity $L$.
@@
We calculate the complexity by abstracting the planning problem for multiple robots by moving each of them sequentially using the $r_{s}$ order of the Lie bracket. Initially, $r_{k}=r_{s}$, there is only one robot contained in each Lie bracket. The complexity in this case is linear in the number of robots since moving one robot to the goal takes constant time. Secondly, when $r_{k}=r_{s}-c$, each Lie bracket contains multiple robots. Moving only one robot is equivalent to moving in the $c$-th order Lie bracket given $r_{k}$ order Lie brackets as controls. Thus, the complexity is calculated as $o(\mathcal{M})=o(nL^{c})=o(nL^{r_{s}-r_{k}})$. In the real world, planning for multiple robots occurs simultaneously, so we consider this complexity as a loose upper bound. ∎
@@
Continuing with the example of 6 robots with 4 groups, consider the motion planning problem $\mathcal{M}(n=4,r_{k}=3)$ where the minimum order to move each robot is $r_{s}=r_{k}=3$. Further, we select six primitives at the highest order Lie brackets that move each robot individually, as shown in Table.2. The problem $\mathcal{M}(4,3)$ reduces to rotating each robot to the goal and then moving it using its corresponding primitive. Since this plan for each robot takes constant time, the complexity of the motion planning problem $\mathcal{M}(4,3)$ is linear in the number of robots.
@@
For the motion planning problem $\mathcal{M}(n=4,r_{k}=2)$, choosing the primitive at the first order Lie brackets. Each primitive moves two robots simultaneously; for example $[h_{2},g_{1}]$ moves robot 4 and 5 sideways and $[h_{2},g_{3}]$ moves robot 1 and 5 sideways. To move each robot individually, the sequence of primitives must mimic the higher-order Lie brackets. For instance, robot 1 is moved in the direction of Lie bracket $[h_{1},[h_{2},g_{3}]]$ using a sequence of primitive chosen from $[h_{2},g_{3}]$ and $h_{1}$. Thus, moving one robot requires a complexity of $o(L^{3-2})=o(L)$ and the overall complexity for $\mathcal{M}(4,2)$ is $o(nL)$.
Lie brackets Robot
$[h_{1},[h_{2},g_{3}]]$ $-c_{1},-s_{1}$
$[h_{1},[h_{3},g_{2}]]$ $-c_{2},-s_{2}$
$[h_{1},[h_{3},g_{2}]]-[h_{2},[h_{2},g_{1}]]$ $-c_{3},-s_{3}$
$[h_{3},[h_{2},g_{1}]]$ $-c_{4},-s_{4}$
$[h_{1},[h_{2},g_{3}]]-[h_{1},[h_{1},g_{3}]]$ $-c_{5},-s_{5}$
$[h_{3},[h_{2},g_{1}]]-[h_{2},[h_{2},g_{1}]]$ $-c_{6},-s_{6}$
Table 2: The min order Lie brackets move along each individual robot. In the right column, only then nonzero elements of each Lie bracket the vector field are shown; $c_{i}$ and $s_{i}$ correspond to $\cos{\theta_{i}}$ and $\sin{\theta_{i}}$ respectively, where $\theta_{i}$ is the orientation of robot $i$.
@@
In short, motion planning is simplified if each robot moves sequentially; however, this approach sacrifices path optimality. The complexity of motion planning increases exponentially as robots get coupled in groups, but this trade-off for parallelism results in more optimal paths.
5.2 Instantiations of motion planning problem
The motion planning problem $\mathcal{M}(n,r_{k})$ addresses a class of problems involving $n$ robots with primitives at most $r_{k}$ order. Problems within this class are not unique but share the same complexity. In this section, we will discuss the instantiation of $\mathcal{M}(n,r_{k})$.

### v2

@@ 4.2 Control of Robot Orientation
We have shown that each group $G_{i}$ induces a vector field $g_{i}$ that only translates the robots in the group and a vector field $h_{i}$ that just rotates the robots that are not in $G_{i}$. What we show next is that we can independently control the orientation of at least 3 robots:
@@
The orientation of any 3 robots can be controlled to any desired orientation $(\theta_{1},\theta_{2},\theta_{3})$. Moreover, at most $m$ robots can be controlled to any desired orientation $(\theta_{1},\ldots,\theta_{m})$.
Remark 4
The robots that are not directly controlled will rotate in place to some unspecified orientation.
Proof
Consider the rotation vector fields $h_{1},\ldots,h_{m-1}$ in (6) and the additional original rotation-only vector field $f_{m}$. All of these vector fields can rotate the robots without translating them. The proposition can then be proved by observing that, using these vector fields, the orientations of the robots linearly depend on the arguments (rotation angles) of these vector fields. For example, for the 4-group case in Table 1, we get the equation:
$\begin{bmatrix}1&1&0&1\\ 1&0&1&1\\ 1&0&0&1\\ 0&0&1&1\\ 0&1&0&1\\ 0&1&1&1\end{bmatrix}\cdot\begin{bmatrix}u_{1}&u_{2}&u_{3}&u_{4}\end{bmatrix}^{T}=q[2n+1:3n]-q[2n+1:3n](0)$ (7)
@@
It is evident that for $m$ groups, the rank of the matrix in the equation above is at most $m$. This implies that at most $m$ of the $n$ equations can be solved exactly for the desired inputs $u_{i}$, indicating that the orientation can be controlled for at most $m$ robots.
Next, we illustrate that the orientation of any 3 robots can be controlled. To see this, note that if one considers the first $m-1$ columns of the matrix, any two rows are distinct (since no two robots belong to the same set of groups), which means that at least two of them are linearly independent. Adding the last column with all ones, it is clear that any additional row will be linearly independent from these two rows, which means that the matrix will have a rank of at least 3 and that any 3 of the $n$ equations can be solved exactly for the desired inputs $u_{i}$. ∎
@@ 4.3 Small-Time Local Controllability (STLC)
After we have identified the control vector fields in Eq. (6), we can show that the swarm is STLC in positions. Traditionally, we can do this by analyzing the closure of $\mathcal{F}$ under the Lie bracket operation. However, we will show that we can independently translate any robot in an arbitrary direction without moving any other robot. This shows that the system is STLC in positions.
Suppose we want to independently translate some robot $k\in\{1,\ldots,n\}$. Let $G_{\hat{k}}$ be any group that contains the robot $k$. Let $\operatorname{Rot}_{i_{1},i_{2},i_{3}}(\theta_{1},\theta_{2},\theta_{3})$ correspond to setting the orientation of robots $i_{1},i_{2},i_{3}$ to $\theta_{1},\theta_{2},\theta_{3}$ as described above. Also, assume that the members of the group $G_{\hat{k}}$ are robots $l_{1},\ldots,l_{j},k$.
@@
The control sequence $P_{1}(d)=(g_{\hat{k}}(d),\operatorname{Rot}_{l_{1},l_{2},k}(\pi,\pi,0),g_{\hat{k}}(d))$ only translates robots $l_{3},\ldots,l_{j},k$.
@@
Under the control sequence above, the first $2$ robots in $G_{\hat{}}$ will travel back and forth while the robot $k$ will undergo translation for $2d$. The rest of the robots in the group will undergo some unspecified motion.
Once we have the control sequence $P_{1}$, it is easy to see that a control sequence $P_{2}=(P_{1}(d),\operatorname{Rot}_{l_{3},l_{4},k}(\pi-\Delta\theta_{l_{3}},\pi-\Delta\theta_{l_{4}},0),P_{1}(d))$ will only translate robots $l_{5},\ldots,l_{j},k$, where $\Delta\theta_{l_{3}},\Delta\theta_{l_{4}}$ are rotations of robots $l_{3},l_{4}$ caused by $P_{1}(d)$. Repeating this process at most $l_{j}/2$ times allows us to eliminate the translation of all robots $l_{1},\ldots,l_{j}$ except for the robot $k$. Of course, we can always initially rotate the robot $k$ so that the final translation is in the desired direction. This shows that each robot can be translated independently and that the system is STLC in positions.∎
@@ 5 Motion Planning
After we have shown that the MicroStressBot swarm under group-control is STLC, we consider the motion planning problem for the swarm. Given our proof of STLC, the problem is not how to move the swarm to a particular configuration, but how to move it efficiently. In other words, ideally, we would like the robots to move to their desired configuration (only the position matters) in parallel and not one by one. However, planning such parallel motion directly is complex due to the high dimension of the configuration space and the fact that we have a single control input. In this paper, we argue that effective motion planning for global group-control in high-dimensional state spaces should involve the design of motion primitives. In particular, motion primitives will be associated with subgroups, trading off the complexity of implementing additional groups in hardware (using PFSMs) with the complexity of control.
The idea of motion primitives is to divide a robot group into smaller subgroups. These primitives simplify the motion planning task by reducing the coupling between robots in the group. We present primitives inspired by Lie bracket motions, following the intuition that, in general, the higher the order of the Lie bracket, the fewer robots are moved by it. It is important to note that, in this paper, these subgroups are logical constructs as opposed to $\log(n+2)+1$ physical groups. However, these logical groups could be implemented in hardware, increasing the number of physical groups and simplifying motion planning and control for the swarm. Of course, this would come with increased fabrication complexity and cost.
In the following section, we are interested in the complexity of the motion planning problem in terms of computation time, path length, and execution time. We investigate how the motion planning problem can be simplified by using motion primitives through both a theoretical complexity bound and simulated instantiations. A pivotal aspect is that a primitive corresponding to a smaller subgroup will generally involve the recursive application of primitives associated with larger subgroups, as observed in the proof of Proposition 4. Because of this, such primitives will involve many back-and-forth motions of the robots, resulting in longer path lengths and execution times. However, motion plans that involve such primitives will be simpler, which means a shorter computation time. On the other hand, when the primitives used for motion planning involve larger subgroups (not much back-and-forth), the motion of the robots is coupled, which increases the computation time for motion planning but minimizes back-and-forth motions and reduces path length and execution time.
@@ 5.1 Motion Planning Approximation Scheme
Unlike physical groups (encoded through PFSMs), a subgroup and the associated motion primitive are a logical construct. A subgroup contains fewer robots than physical groups, and only those robots move forward when a subgroup is activated. This abstraction will allow us to formulate the motion planning problem in layers of increasing difficulty. Physically, the motion induced by a given subgroup $p$ corresponds to a particular Lie bracket and can thus be physically realized. We will refer to such motion as a motion primitive associated with the subgroup $p$.
@@
[Primitive Order] The order of a primitive corresponding to the subgroup $p$ is the order of the Lie bracket used to realize it. The order of the Lie bracket is the number of generators from the set of vector fields $\mathcal{F}$ in Eq. (6) in it.
In general, the higher the order of a Lie bracket, the fewer robots will move by that bracket. However, there is no one-to-one correspondence between the order of the Lie bracket and the number of robots it moves, as this also depends on the vector fields involved in the Lie bracket. Fundamentally, the order of the Lie bracket indicates the difficulty of its implementation; higher order is more difficult to implement. On the other hand, the number of robots affected by a Lie bracket reflects the complexity of motion planning; the more robots move simultaneously, the more difficult it is to move each of them to the desired configuration.
@@
Motion Planning Problem Abstraction The motion planning problem for $n$ robots of level $r$, denoted as $\mathcal{M}_{q_{s}}^{q_{g}}(n,r)$, corresponds to the problem of finding a collision-free trajectory (sequence of motion primitives) for $n$ robots controlled by a global field from some initial state $q_{s}$ to a goal state $q_{g}$, where the motion primitives that are used have the primitive order at most $r$.
@@
Consider the motion planning problem $\mathcal{M}_{q_{s}}^{q_{g}}(n,k)$ for $n$ robots and $1\leq k\leq k_{max}$, where $k_{max}$ is the primitive order required for STLC. If $k_{1}<k_{2}\leq k_{max}$, the complexity of $\mathcal{M}_{s}^{g}(n,k_{2})$ is less than $\mathcal{M}_{q_{s}}^{q_{g}}(n,k_{1})$. When $k=k_{max}$, the complexity of $\mathcal{M}_{q_{s}}^{q_{g}}(n,k_{max})$ solved by RRT [14] is $O(n\cdot c^{n})$. For any other $k$, the complexity of $\mathcal{M}_{q_{s}}^{q_{g}}(n,k)$ solved by RRT is $O(n\cdot c^{n}\cdot L^{k_{max}-k})$, where $L$ is the complexity of the first-order Lie bracket and $c$ is a constant depending on the environment size and resolution.
Remark 5
By observing that the primitives used in the proof of Theorem 4 are analogous to a Lie bracket operation, it can be shown that $k_{max}<=m$. The formal proof relies on the special properties of the vector fields in $\mathcal{F}$ (e.g., $[h_{i},h_{j}]=[g_{i},g_{j}]=[h_{i},g_{i}]=0$ for any $i,j<m$) and is beyond the scope of this paper.
@@
For $k=k_{max}$, the system is fully actuated due to STLC, allowing motion in all directions of the configuration space. For position control of $n$ robots, the dimension of the configuration space is $d=2n$. As the complexity of representing the space is exponential in $d$, the number of samples $n_{s}$ required for RRT scales as $O(c^{n})$, resulting in the complexity $O(n_{s}\log n_{s})=O(c^{n}\log(c^{n}))=O(n\cdot c^{n})$ for the RRT solver.
When $k<k_{max}$, the system is underactuated, meaning that motion in certain directions cannot be directly achieved. Instead, these motions are approximated through paths composed of Lie brackets of control vector fields. The Lie bracket is implemented by alternating the motion back and forth along two vector fields. If $L$ is the complexity of this maneuver for the first-order Lie bracket, higher-order Lie brackets correspond to nested applications of an analogous maneuver. Given a Lie bracket of order $k$, the motion in a given direction must ultimately be implemented through the Lie brackets of the order $k_{max}$. The complexity of this implementation is $L^{k_{max}-k}$. Therefore, the overall complexity of the RRT algorithm becomes $O(n_{s}L^{k_{max}-k}\log n_{s})=O(n\cdot c^{n}\cdot L^{k_{max}-k})$.
As $k$ approaches $k_{max}$, the complexity reduces since the system becomes more actuated, requiring fewer nested Lie bracket computations. Consequently, the complexity of $\mathcal{M}_{q_{s}}^{q_{g}}(n,k_{2})$ is less than $\mathcal{M}_{q_{s}}^{q_{g}}(n,k_{1})$ for $k_{1}<k_{2}\leq k_{max}$. ∎
@@
Continuing with the example of 6 robots and 4 groups, consider the motion planning problem $\mathcal{M}(6,2)$. In this case, the minimum order to move each robot is $r_{s}=2$ (see Table 2). We can select six primitives involving these second-order Lie brackets that move each robot individually. The problem $\mathcal{M}(6,2)$ then reduces to rotating each robot toward the goal and moving it using its corresponding primitive.
@@
For the motion planning problem $\mathcal{M}(6,1)$, the original vector fields from $\mathcal{F}$ are chosen as the primitives. Each primitive moves three robots simultaneously. For example, $g_{1}$ moves robots 4, 5, and 6, and $g_{3}$ moves robots 1, 3, and 5. To move each robot to its desired configuration using these primitives requires the application of the Lie bracket $[h_{1}+h_{2}-f_{4},g_{3}]$ utilizing a sequence of primitives $g_{3}$ and $h_{1}+h_{2}f_{4}$.
Lie brackets Robots
$[h_{2},g_{1}]$ 4, 5
$[h_{3},g_{1}]$ 4, 6
$[h_{1},g_{2}]$ 2, 3
$[h_{3},g_{2}]$ 2, 6
$[h_{1},g_{3}]$ 1, 3
$[h_{2},g_{3}]$ 1, 5
Lie brackets Robots
$[h_{1}+h_{2}-f_{4},g_{3}]$ 1
$[h_{1}+h_{3}-f_{4},g_{2}]$ 2
$[f_{4}-h_{3},g_{2}]$ 3
$[h_{2}+h_{3}-f_{4},g_{1}]$ 4
$[f_{4}-h_{3},g_{1}]$ 5
$[f_{4}-h_{2},g_{1}]$ 6
Table 2: Lie brackets for $m=4$ and robots that are moved by them.
@@
In short, motion planning is simplified if fewer robots are moved by a primitive; however, this approach compromises path efficiency. The computational complexity of motion planning increases exponentially as primitive order decreases and the motion of more robots is coupled; however, the robots move in parallel, so the paths are more efficient.
6 Instantiations of the motion planning problem
The motion planning problem $\mathcal{M}(n,k)$ involves moving $n$ robots with primitives of order $k$. This section investigates some instantiations of the motion planning problem $\mathcal{M}(n,k)$.